As investment with a 3year life and a cost of #120000 generates revenue of #25000 in year 1 #45000 in year 2 and #65000 in year 3. If the discount rate is 8%

(i) what is the NPV of the investment.

(ii) state whether the investment should be accepted or not and why.

(iii) using 12% as your second discount rate, to solve the remaining part of the question under IRR.


1
Expert's answer
2021-08-05T13:58:52-0400

Solution


"YEAR_ 1= \frac{25000}{1.08}=23148.14"


"YEAR _2=\frac{45000}{1.08^2}=38580.25"


"YEAR_ 3=\frac{65000}{1.08^3}=51599.10"



"NPV= -120000+23148.15+38580.25+51599.1"


NPV1= -6672.5


Ii.) Negative NPV, investment should not be accepted


12% discount

PV= 120000


"YEAR_1=\frac{25000}{1.12}=22321.42"



"YEAR_2=\frac{45000}{1.12^2}=35873.72"



"YEAR_3=\frac{65000}{1.12^3}=46265.72"



"NPV_2=\u2212120000+22321.43+35873.72+46265.72\n\n= -15539.13"

t=3

"IRR=\sum\frac{C_t}{(1+r)^t}-C_o"

t=1

t= number of years


Co = initial investments


t= Number or periods


r= Discount rate


"IRR= 6.469" %